OpenStudy (df001):

Help proof: 1-sin(x)/cos(x) = cos(x)/1+sin(x)

OpenStudy (df001):

\[\frac{ 1-\sin x }{ \cos x } = \frac{ \cos x }{ 1+\sin x }\]

OpenStudy (df001):

i need help on this last problem before i sleep pls help me :(

OpenStudy (df001):

it's so confusing i don't know where to start.. all I can see is 1-sin x = cos if i am correct

OpenStudy (df001):

but cos^2 = 1-sin^2x.. i dont even know if that works

OpenStudy (df001):

hello pls help me

OpenStudy (df001):

hi

OpenStudy (df001):

i am very stuck on this problem.. im not sure if cosx can cancel out

OpenStudy (df001):

hello pls help me

OpenStudy (df001):

my brian hurts :[

Directrix (directrix):

Work on the right side only. Begin by multiplying the right side by (1-sin x) / (1 - sin x}